OpenStudy (anonymous):

Evaluate: The limit as theta approaches 0 of sin theta cot 2 theta

OpenStudy (anonymous):

\[\lim_{x\rightarrow 0}\frac{\sin(x)}{\tan(2x)}\] now l'hopital should give it can you use that one?

OpenStudy (anonymous):

Were not suppose to know about that rule yet were supposed to do it the other way algebraically

OpenStudy (zarkon):

\[\lim_{x\rightarrow 0}\frac{\sin(x)}{\tan(2x)}\] \[=\lim_{x\rightarrow 0}\frac{1}{2}\frac{2x}{\tan(2x)}\frac{\sin(x)}{x}\]

OpenStudy (anonymous):

zarkon! hardly recognized you

OpenStudy (anonymous):

So the answer is 1/2?

OpenStudy (zarkon):

I've reverted back to a cartoon form yes...the answer is 1/2

OpenStudy (anonymous):

Thank you!